OpenStudy (anonymous):

Explain why the logarithm of 1 with base b is 0.

OpenStudy (lgbasallote):

because anything raised to 0 is 1

OpenStudy (anonymous):

x = log(base b) of y is equivalent to y = b^x So if y = 1, we have 1 = b^x which implies that x = 0 since b is not 0.

Parth (parthkohli):

When b \(\ne\) 0.

Parth (parthkohli):

@Rdx you're fast at copy/pasting

OpenStudy (anonymous):

yes its from yahoo answers i know cause i asked the same question to myself too... and searched on google. so instead on writing it i just copied it from their and wrote it, but if it is something against the rule i m sorry . I am new here

OpenStudy (anonymous):

@ParthKohli ^^^^^

OpenStudy (anonymous):

\[b^x=y\iff \log_b(y)=x\] \[b^0=1\iff \log_b(1)=0\] that is why